# non-asymptotic estimates for markov transition matrices with rigorous error bounds

### v1

@@ Abstract.
Then $\mu(u)$ and $p(u,v)$ can be approximated by $\sum_{w\in\Omega}r_{n}(u,w)$ and $r_{n}(u,v)/\sum_{w\in\Omega}r_{n}(u,w)$, respectively. We will use $\bm{R}_{n}=[r_{n}(u,v)]_{u,v\in\Omega}$ to denote the matrix form of MLE. It is shown in [1, 3] that the error of MLE is asymptotically $\mathcal{O}(1/\sqrt{n})$ by the central limiting theorem for Markov chains. However, to the best of our knowledge, there is no quantitative result on the non-asymptotic behavior of MLE in the current literature.
Our first contribution is to establish non-asymptotic error bounds of MLE. To achieve this, we interpret MLE as the Markov empirical mean of the indicator function $\mathbb{1}_{(u,v)}(\cdot,\cdot)$ corresponding to some transition matrix $\bm{P}_{2}$ defined on the length-2 path space $\Omega^{2}$. We then study the spectrum of $\bm{P}_{2}$ and apply concentration inequalities for Markov chains to prove the following entry-wise tail bound for the error of MLE.
@@
$\mathbb{P}\left\{|r_{n}(u,v)-\mu(u)p(u,v)|\geq t\right\}\leq C_{1}\|\nu/\mu\|_{\infty}\exp\left(-C_{2}\eta_{s}(\bm{P})nt^{2}\right),$ (1.2)
where $C_{1}$, $C_{2}$ are absolute constants, $\nu$ is the initial distribution of $u_{1}$, and $\eta_{s}(\bm{P})$ is the symmetric spectral gap of $\bm{P}$.
There has been extensive research on concentration/deviation inequalities for Markov chains [11, 12, 13, 17, 6, 10]. Most of these results presented in these studies typically require the transition matrix to satisfy certain spectral conditions, which $\bm{P}_{2}$ fails to meet. We also notice that the proofs of concentration inequalities for general Markov chains given in [6, 10] seem to have some gap. Therefore, we choose to adopt the result in [8], that can be directly applied to $\bm{P}_{2}$.
In numerous applications, the Markov chain considered is reversible [7], i.e. the transition matrix satisfies the detailed balance condition. In this case, it is important that the estimator preserves the reversibility of $\bm{P}$  [2, 18, 14]. Similarly, the theories proposed in the literature do not provide any non-asymptotic convergence rate guarantee for those estimators.
@@
Non-asymptotic expected error bounds are sometimes more favorable in applications. A straightforward way to bound the error expectation is by integrating the corresponding tail bound. However, for MLE and SCE, the expectation bounds derived in this way both depend on the dimension $d$, which is inconsistent with our numerical observations. Instead, we derive dimension-free expectation bounds of both estimators by utilizing the special structure $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$.
@@
$\operatorname{\mathbb{E}}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}\right]=\mathcal{O}\left(\frac{1}{\sqrt{n\eta_{s}(\bm{P})}}\right),$ (1.5)
where $\eta_{s}(\bm{P})$ is the symmetric spectral gap of $\bm{P}$. Moreover, when $\bm{P}$ is reversible,
@@ 2.1.1. Notation
Let $\Omega$ be a Polish space equipped with a probability measure $\mu$. Define $\mathbb{E}_{\mu}$ the expectation with respect to the measure $\mu$. When applied to a random matrix, $\mathbb{E}_{\mu}$ computes the entry-wise expectation. In this paper, the number of states $|\Omega|$ is assumed to be finite and is always denoted by $d$. Any real-valued function on $\Omega$ can be seen as a finite-dimensional real vector and vice versa.
@@ 2.1.2. Foundation of Markov chains
For a function $f:\Omega\rightarrow\mathbb{R}$, the action of $\bm{P}$ as a linear operator on $f$ is defined as
@@
2.1.3. Absolute spectral gap and symmetric spectral gap
Throughout this paper, we will assume that the transition matrix $\bm{P}$ is irreducible, and thus its invariant distribution $\mu$ is unique. Let $L_{2,\mu}$ be the Hilbert space of all real-valued functions on $\Omega$ endowed with the inner product
@@ 2.1.3. Absolute spectral gap and symmetric spectral gap
and $L_{2,\mu}^{0}=\{f\in L_{2,\mu}:\operatorname{\mathbb{E}}_{\mu}[f]=0\}$ be the mean-zero subspace of $L_{2,\mu}$. We will use $\|\cdot\|_{\mu}$ to denote both the $L_{2,\mu}$ norm of a vector and the $L_{2,\mu}$ operator norm of a matrix. Write $\bm{\Pi}=\bm{1}\mu^{\mathrm{T}}$, the projection onto $\bm{1}$ under the $L_{2,\mu}$ inner product. One can verify that $\bm{\Pi}\bm{P}=\bm{P}\bm{\Pi}=\bm{\Pi}$, and therefore
$\bm{P}^{k}-\bm{\Pi}=(\bm{P}-\bm{\Pi})^{k}=(\bm{P}-\bm{\Pi})^{k}(\bm{\mathbf{I}}-\bm{\Pi}),\quad\text{for any}\,\ k\in\mathbb{N}_{+}.$ (2.6)
We say the transition matrix $\bm{P}$ admits a nonzero absolute spectral gap $\eta_{a}(\bm{P})=1-\lambda_{a}(\bm{P})$ if
$\lambda_{a}(\bm{P}):=\sup_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\|\bm{P}h\|_{\mu}}{\|h\|_{\mu}}=\|\bm{P}-\bm{\Pi}\|_{\mu}<1.$ (2.7)
Let $\bm{P}^{*}$ be the adjoint of $\bm{P}$ with respect to the $L_{2,\mu}$ inner product and $\bm{A}:=(\bm{P}+\bm{P}^{*})/2$ be the additive symmetrization of $\bm{P}$. We say $\bm{P}$ admits a nonzero symmetric spectral gap $\eta_{s}(\bm{P})=1-\lambda_{s}(\bm{P})$ if
$\lambda_{s}(\bm{P}):=\sup_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\left\langle h,\bm{P}h\right\rangle_{\mu}}{\|h\|_{\mu}^{2}}=\sup_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\left\langle h,\bm{A}h\right\rangle_{\mu}}{\|h\|_{\mu}^{2}}<1.$ (2.8)
By the irreducibility of $\bm{P}$ and the Perron-Frobenius theorem, we know $\mu(u)\neq 0$ for any $u\in\Omega$, and thus $\bm{D}_{\mu}$ is invertible. It is not hard to find that $\lambda_{a}(\bm{P})$ is the second largest singular value of the matrix $\bm{D}_{\mu}^{1/2}\bm{P}\bm{D}_{\mu}^{-1/2}$, which is also the largest singular value of $\bm{D}_{\mu}^{1/2}(\bm{P}-\bm{\Pi})\bm{D}_{\mu}^{-1/2}$, and $\lambda_{s}(\bm{P})$ is the second largest eigenvalue of $(\bm{D}_{\mu}^{1/2}\bm{P}\bm{D}_{\mu}^{-1/2}+\bm{D}_{\mu}^{-1/2}\bm{P}^{\mathrm{T}}\bm{D}_{\mu}^{1/2})/2$. By Jensen’s inequality,
Hence $\lambda_{a}(\bm{P})\leq 1$. Moreover, a direct application of the Cauchy-Schwarz inequality implies $|\lambda_{s}(\bm{P})|\leq\lambda_{a}(\bm{P})$, and therefore $\eta_{s}(\bm{P})\geq\eta_{a}(\bm{P})$. One should realize that $\bm{A}$ is also a transition matrix (the sum of each row is $1$), and it is irreducible since $\bm{P}$ is irreducible. Therefore, we always have $\lambda_{s}(\bm{P})<1$. However, irreducibility does not guarantee that $\lambda_{a}(\bm{P})<1$. In fact, in Section 3 we will construct a family of irreducible transition matrices whose absolute spectral gaps are zero.
In particular, when $\bm{P}$ is reversible, $\lambda_{s}(\bm{P})$ is the second largest eigenvalue of $\bm{P}$ and $\eta_{s}(\bm{P})=1-\lambda_{s}(\bm{P})$ is the gap between the largest two eigenvalues of $\bm{P}$. In this case we will omit the subscript $s$ to use $\lambda(\bm{P})$ for the second largest eigenvalue of $\bm{P}$ and $\eta(\bm{P})=1-\lambda(\bm{P})$ for the spectral gap of $\bm{P}$.
For technical convenience, in the rest of this paper we will assume $\bm{P}$ is a dense matrix, i.e. $p(u,v)\neq 0$ for any $u,v\in\Omega$. Note that all spectral properties discussed above are well-posed under small perturbation on $\bm{P}$, and hence all arguments and conclusions in this paper can be readily generalized to non-dense cases by the limiting argument.
@@ 2.2. Main results
    $\mathbb{P}\left\{\left|r_{n}(u,v)-\mu(u)p(u,v)\right|\geq t\right\}\leq 2\|\nu/\mu\|_{\infty}\exp\left(\frac{-n\eta_{s}(\bm{P})t^{2}}{8\sqrt{\left(2+5\eta_{s}(\bm{P})/2\sqrt{2}\right)^{2}+t^{2}}}\right).$ (2.11)
@@
    $\operatorname{\mathbb{E}}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\right]\leq\|\nu/\mu\|_{\infty}\min\left\{\frac{4+\eta_{s}(\bm{P})}{n\eta_{s}(\bm{P})},\frac{2+\eta_{a}(\bm{P})}{n\eta_{a}(\bm{P})}\right\}.$ (2.12)
@@
When $\bm{P}$ is reversible, the Markov chain constructed in Theorem 2.1 is nonreversible in general. In order to utilize the reversibility of $\bm{P}$, we can symmetrize $\Omega^{2}$ by viewing $(u_{1},u_{2})$ and $(u_{2},u_{1})$ as the same element to define a symmetric length-2 path space $\tilde{\Omega}^{2}$ and construct a reversible transition matrix $\tilde{\bm{P}}_{2}$ on $\tilde{\Omega}^{2}$. When we have access to an online oracle of $\bm{P}$ that takes an input $u\in\Omega$ and outputs a state $v\in\Omega$ with probability $p(u,v)$, we propose a Symmetric Counting Estimation(SCE) algorithm that generates a sequence $\{\tilde{\bm{u}}_{k}\}_{k\geq 1}\subset\tilde{\Omega}^{2}$ according to $\tilde{\bm{P}}_{2}$ and approximates $\bm{D}_{\mu}\bm{P}$ using the symmetric matrix empirical mean

### v2

@@ Abstract.
Then $\mu(u)$ and $p(u,v)$ can be approximated by $\sum_{w\in\Omega}r_{n}(u,w)$ and $r_{n}(u,v)/\sum_{w\in\Omega}r_{n}(u,w)$, respectively. We will use $\bm{R}_{n}=[r_{n}(u,v)]_{u,v\in\Omega}$ to denote the matrix form of MLE. It is shown in [1, 3] that the error of MLE is asymptotically $\mathcal{O}(1/\sqrt{n})$ by the central limiting theorem for Markov chains. In contrast, our goal is to establish non-asymptotic results that are valid for any $n\in\mathbb{N}^{+}$.
Our first contribution is to establish non-asymptotic error bounds of MLE in terms of the iterated Poincaré gap (IP gap) of $\bm{P}$, which was first introduced in [5] as a generalization of the spectral gap for non-reversible Markov chains. We call it the iterated Poincaré gap gap, following the notations in [10], to emphasize its definition by an iterated Poincaré gap inequality (see Subsection 2.1 below). To achieve this, we interpret MLE as the Markov empirical mean of the indicator function $\mathbb{1}_{(u,v)}(\cdot,\cdot)$ corresponding to some transition matrix $\bm{P}_{2}$ defined on the length-2 path space $\Omega^{2}$. We then establish the relation between the spectral gaps of $\bm{P}_{2}$ and $\bm{P}$, and apply concentration inequalities for Markov chains to derive the following entry-wise tail bound for the error of MLE.
@@
$\mathbb{P}\left\{|r_{n}(u,v)-\mu(u)p(u,v)|\geq t\right\}\leq C_{1}\|\nu/\mu\|_{\infty}\exp\left(-C_{2}\eta_{p}(\bm{P})nt^{2}\right),$ (1.2)
where $C_{1}$, $C_{2}$ are absolute constants, $\nu$ is the initial distribution of the Markov sequence, and $\eta_{p}(\bm{P})$ is the IP gap of $\bm{P}$.
There have been many related works that also focus on learning and testing Markov transition matrices [11, 23, 4, 25]. Most relevantly, [23] controls the probabilistic sample complexity of MLE at the matrix level (in terms of the total variation norm). We will discuss their results in detail in Section 2. We want to emphasize that, compared with prior work, the novelty of our result lies in that it only requires a nonzero IP gap of $\bm{P}$ (which proves to be more robust than other generalized spectral gaps for non-reversible Markov chains; see [5, 10]), owing to a recently established Bernstein-type inequality [10] by the same authors. We will show that $\eta_{p}(\bm{P})>0$ holds for any irreducible transition matrix $\bm{P}$. Hence, Theorem 1.1 applies to all Markov chains with irreducible transition matrices, highlighting the robustness of our result.
In numerous applications, the Markov chain considered is reversible [9], i.e. the transition matrix satisfies the detailed balance condition. In this case, it is important that the estimator preserves the reversibility of $\bm{P}$  [2, 21, 17]. However, the theories proposed in the literature do not provide any non-asymptotic convergence rate guarantee for those estimators.
@@
Non-asymptotic expected error bounds are sometimes more favorable in applications. A simple way to bound the error expectation is by integrating the corresponding tail bound, which can be crude since the tail bound is in general not sharp for all scales of deviations. For example, if we integrate the Bernstein-type tail bounds above to control the expected errors for MLE and SCE, the resulting bounds will typically depend on the dimension $d$, which is in fact a technical overestimation. Instead, we derive dimension-free expectation bounds of both estimators via a direct calculation by utilizing the special structure $\bm{P}_{2}$ and $\tilde{\bm{P}}_{2}$.
@@
$\operatorname{\mathbb{E}}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}\right]=\mathcal{O}\left(\frac{1}{\sqrt{n\eta_{p}(\bm{P})}}\right),$ (1.5)
where $\eta_{p}(\bm{P})$ is the IP gap of $\bm{P}$. Moreover, when $\bm{P}$ is reversible,
@@ 2.1.1. Notation
Let $\Omega$ be a finite state space equipped with a probability measure $\mu$, the number of states $|\Omega|$ is denoted by $d$. Any real-valued function on $\Omega$ can be seen as a finite-dimensional real vector and vice versa. Define $\mathbb{E}_{\mu}$ the expectation with respect to the measure $\mu$. When applied to a random matrix, $\mathbb{E}_{\mu}$ computes the entry-wise expectation.
@@ 2.1.2. Foundation of Markov chains
Throughout this paper, we will assume that the transition matrix $\bm{P}$ is irreducible, and thus its invariant distribution $\mu$ is unique. For a function $f:\Omega\rightarrow\mathbb{R}$, the action of $\bm{P}$ as a linear operator on $f$ is defined as
@@
Let $L_{2,\mu}$ be the Hilbert space of all real-valued functions on $\Omega$ endowed with the inner product
@@
and $L_{2,\mu}^{0}=\{f\in L_{2,\mu}:\operatorname{\mathbb{E}}_{\mu}[f]=0\}$ be the mean-zero subspace of $L_{2,\mu}$. We will use $\|\cdot\|_{\mu}$ to denote both the $L_{2,\mu}$ norm of a vector and the $L_{2,\mu}$ operator norm of a matrix.
In the reversible case, $\bm{P}$ is self-adjoint under the $L_{2,\mu}$ inner product, and the spectral gap of the transition matrix $\bm{P}$ is given by $\eta(\bm{P})=:1-\lambda(\bm{P})$, where
$\lambda(\bm{P}):=\sup_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\langle h,\bm{P}h\rangle_{\mu}}{\|h\|_{\mu}^{2}}.$ (2.6)
In other words, $\eta(\bm{P})$ is the gap between 1 and the second largest eigenvalue of $\bm{P}$.
When $\bm{P}$ is non-reversible, we define the the iterated Poincaré gap (IP gap) of $\bm{P}$ as
$\eta_{p}(\bm{P}):=\inf_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\|(\bm{\mathbf{I}}-\bm{P})h\|_{\mu}}{\|h\|_{\mu}}.$
This definition coincides with that of the non-reversible spectral gap first formally introduced by Chatterjee in [5], where he used this quantity to establish ergodicity for non-reversible Markov chains, generalizing similar results for reversible Markov chains based on the usual spectral gap. The reader may thus also refer to the IP gap as the Chatterjee gap or the non-reversible spectral gap and can find a more comprehensive introduction to the topic in [5]. This quantity was independently developed by [10] to establish Bernstein-type inequalities for Markov chains. In this paper, we choose to adopt the notation in [10] and call this quantity the iterated Poincaré gap, since it is related to the iterated Poincaré inequality:
$\operatorname{Var}_{\mu}[h]\leq\eta_{p}^{-2}\mathbb{E}_{\mu}\left[((\bm{\mathbf{I}}-\bm{P})h)^{2}\right],\quad\text{ for any }h\in L_{2,\mu}.$
@@ 2.2. Main results
    $\mathbb{P}\left\{\left|r_{n}(u,v)-\mu(u)p(u,v)\right|\geq t\right\}\leq 2\|\nu/\mu\|_{\infty}\exp\left(\frac{-n\,\gamma(\bm{P})\,t^{2}}{4\sqrt{\left(2+6\gamma(\bm{P})\right)^{2}\mu_{2}(u,v)+t^{2}}}\right),$ (2.8)
    where $\mu_{2}(u,v)=\mu(u)p(u,v)$, $\gamma(\bm{P})=\eta_{p}(\bm{P})/(1+\eta_{p}(\bm{P}))$.
@@
    $\operatorname{\mathbb{E}}\left[\left\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\right\|_{\mathrm{F}}^{2}\right]\leq\|\nu/\mu\|_{\infty}\min\left\{\frac{4+\eta_{p}(\bm{P})}{n\eta_{p}(\bm{P})},\frac{2+\eta_{a}(\bm{P})}{n\eta_{a}(\bm{P})}\right\}.$ (2.9)
@@
When $\bm{P}$ is reversible, the Markov chain constructed in Theorem 2.1 is non-reversible in general. In order to utilize the reversibility of $\bm{P}$, we can symmetrize $\Omega^{2}$ by viewing $(u_{1},u_{2})$ and $(u_{2},u_{1})$ as the same element to define a symmetric length-2 path space $\tilde{\Omega}^{2}$ and construct a reversible transition matrix $\tilde{\bm{P}}_{2}$ on $\tilde{\Omega}^{2}$. When we have access to an online oracle of $\bm{P}$ that takes an input $u\in\Omega$ and outputs a state $v\in\Omega$ with probability $p(u,v)$, we propose a Symmetric Counting Estimation(SCE) algorithm that generates a sequence $\{\tilde{\bm{u}}_{k}\}_{k\geq 1}\subset\tilde{\Omega}^{2}$ according to $\tilde{\bm{P}}_{2}$ and approximates $\bm{D}_{\mu}\bm{P}$ using the symmetric matrix empirical mean
@@
2.3. Generalized spectral gaps
In the non-reversible case, various notions of spectral gaps have been proposed as substitutes for the usual spectral gap before [5]. A commonly used quantity is the absolute $L_{2}$-spectral gap (or simply the absolute spectral gap), which is defined as $\eta_{a}(\bm{P})=1-\lambda_{a}(\bm{P})$, where
$\lambda_{a}(\bm{P}):=\sup_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\|\bm{P}h\|_{\mu}}{\|h\|_{\mu}}=\|\bm{P}-\bm{\Pi}\|_{\mu}.$ (2.14)
Here $\bm{\Pi}:=\bm{1}\mu^{\mathrm{T}}$ is the projection onto $\bm{1}$ under the $L_{2,\mu}$ inner product. One can see that $\eta_{a}(\bm{P})$ is the gap between $1$ and the second largest singular value of $\bm{P}$ under the $L_{2,\mu}$ inner product. Alternatively, $\eta_{a}(\bm{P})$ can be understood as the usual spectral gap of the self-adjoint operator $\left(\bm{P}^{*}\bm{P}\right)^{1/2}$, where $\bm{P}^{*}$ is the adjoint of $\bm{P}$ under the $L_{2,\mu}$ inner product. This quantity was widely used in the literature to bound the mixing time [7] and establish concentration inequalities [14, 15] for non-reversible Markov chains.
A generalization of the absolute spectral gap, called the pseudo spectral gap, was first introduced in [20] to establish concentration inequalities for non-reversible Markov chains. It is defined as
$\eta_{ps}(\bm{P}):=\sup_{k\in\mathbb{N}_{+}}\frac{1}{k}\left(1-\lambda_{a}^{2}\left(\bm{P}^{k}\right)\right)=\sup_{k\in\mathbb{N}_{+}}\frac{\eta\left((\bm{P}^{*})^{k}\bm{P}^{k}\right)}{k}.$
Advanced theoretical studies, estimations and applications of the pseudo spectral gap can be found in [22, 24].
Another way to define the spectral gap for a non-reversible Markov chain is by simply using the usual spectral gap of the self-adjoint operator $\bm{A}:=(\bm{P}^{*}+\bm{P})/2$, the additive symmetrization of $\bm{P}$. To be specific, the symmetric spectral gap of $\bm{P}$ is defined as $\eta_{s}(\bm{P})=:1-\lambda_{s}(\bm{P})$, where
$\lambda_{s}(\bm{P}):=\sup_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\langle h,\bm{P}h\rangle_{\mu}}{\|h\|_{\mu}^{2}}=\sup_{h\in L_{2,\mu}^{0},\ h\neq\bm{0}}\frac{\langle h,\bm{A}h\rangle_{\mu}}{\|h\|_{\mu}^{2}}=\lambda(\bm{A}),$
It is shown in [7] that $\eta_{s}(\bm{P})$ can be used to get upper bounds on the mixing time of $\bm{P}$ when $\bm{P}$ is strongly aperiodic. Moreover, the additive reversiblization $\bm{A}$ is usually more user-friendly than the the multiplicative reversiblization $\bm{P}^{*}\bm{P}$. However, except for [8, 13], which appear to contain a technical gap in the proof of a key lemma (see [10] for details), there are no concentration results that depend on $\eta_{s}(\bm{P})$.
Note that in the reversible case, it always holds that $\eta(\bm{P})=\eta_{p}(\bm{P})=\eta_{s}(\bm{P})$, and if the second largest eigenvalue of $\bm{P}$ in absolute value is positive, then it also holds that $\eta_{p}(\bm{P})=\eta_{a}(\bm{P})$. When $\bm{P}$ is non-reversible, one can prove that the relations $\eta_{p}(\bm{P})\geq\eta_{s}(\bm{P})\geq\eta_{a}(\bm{P})$ and $\eta_{p}(\bm{P})\geq\eta_{ps}(\bm{P})/2$ (Lemma A.1) for any non-reversible Markov transition matrix. Moreover, by the Perron–Frobenius theorem, the irreducibility of $\bm{P}$ (which implies the irreducibility of $(\bm{P}^{*}+\bm{P})/2$) guarantees that $\eta_{p}(\bm{P})\geq\eta_{s}(\bm{P})>0$. However, irreducibility does not guarantee that $\eta_{a}(\bm{P})>0$. We can only show that $\lambda_{a}(\bm{P})\leq 1$ (and hence $\eta_{a}(\bm{P})\geq 0$) by Jensen’s inequality:
In fact, there are simple examples where $\eta_{p}(\bm{P})>0$ but the ratios $\eta_{a}(\bm{P})/\eta_{p}(\bm{P})$ and $\eta_{ps}(\bm{P})/\eta_{p}(\bm{P})$ can be arbitrarily small or even zero (see Appendix A).
2.4. Related works
There have been many related works that also focus on learning and testing Markov transition matrices [11, 23, 4, 25]. Most relevantly, [23] controls the probabilistic sample complexity of MLE in terms of the mixing time $t_{\text{mix}}$. More specifically, we say the transition matrix $\bm{P}$ is geometrically ergodic, if there exists $\rho\in(0,1)$ such that for any $u\in\Omega$,
$\|\bm{P}^{n}(u,\cdot)-\mu^{\mathrm{T}}\|_{\text{TV}}\leq C\rho^{n},$
where $\|\cdot\|_{\text{TV}}=\|\cdot\|_{l_{1}}/2$ is the standard total variation norm. The mixing time of $\bm{P}$ is then defined by
$t_{\text{mix}}:=\inf_{n\in\mathbb{N}_{+}}\sup_{\nu}\|\nu^{\mathrm{T}}\bm{P}^{n}-\mu\|_{\text{TV}},$
where the supremum is taken over all distributions over $\Omega$.
It is shown in [23, Theorem 3.3] that if $\bm{P}$ is geometrically ergodic, then for any $\varepsilon,\delta>0$ and any
$n\geq c\frac{t_{\text{mix}}}{\varepsilon^{2}}\max\left\{\|\bm{D}_{\mu}\bm{P}\|_{1/2},\log\left(\frac{\|\nu/\mu\|_{\mu}}{\delta}\right)\right\},$
$\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\|_{\text{TV}}<\epsilon$ holds with probability at least $1-\delta$. Here $c$ is an absolute constant, $\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\|_{\text{TV}}$ is the total variation error between $\bm{R}_{n}$ and $\bm{D}_{\mu}\bm{P}$ (interpreted as distributions over $\Omega^{2}$), and
$\|\bm{D}_{\mu}\bm{P}\|_{1/2}:=\left(\sum\limits_{u,v\in\Omega}(\mu(u)p(u,v))^{1/2}\right)^{2}.$
It is not hard to see that $\|\bm{D}_{\mu}\bm{P}\|_{1/2}$ scales like $d^{2}$ in the worst case. Hence, the resulting sample complexity may depend on the dimension $d$.
[…]
$\mathbb{P}\left\{{|\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\|_{\text{TV}}-\mathbb{E}[\|\bm{R}_{n}-\bm{D}_{\mu}\bm{P}\|_{\text{TV}}]|>\varepsilon/2}\right\}\leq 2\exp\left(-\frac{n\varepsilon^{2}}{ct_{\text{mix}}}\right).$ (2.16)
Finally, the result was extended to the non-stationary case with an extra factor $\|\nu/\mu\|_{\mu}$.
It is worth-noting that [23, Theorem 3.3] only works when the transition matrix $\bm{P}$ is geometrically ergodic. However, by the central limiting theorem for Markov chains, one does not need the geometrically ergodicity of $\bm{P}$ to guarantee the convergence of the Markov empirical mean to its space average. In comparison, our Theorem 2.1(2) establishes probabilistic error bound for MLE at the entry-wise level, which can be applied to any irreducible Markov chains (not necessarily geometrically ergodic). Note that $r_{n}(u,v)$ can be interpreted as the Markov empirical mean of the indicator function $\mathbb{1}_{(u,v)}(\cdot,\cdot)$ corresponding to $\bm{P}_{2}$. Therefore, a recently established Bernstein-type inequality for Markov sums [10], under milder assumptions than [20, Corollary 2.10], can be applied to $\bm{P}_{2}$ to derive that
$\mathbb{P}\left\{{|r_{n}(u,v)-\mu(u)p(u,v)|>t}\right\}\lesssim\exp\left(-cn\eta_{p}(\bm{P}_{2})t^{2}\right).$
We then establish Theorem 2.1(2) by further showing that $\eta_{p}(\bm{P}_{2})\geq\eta_{p}(\bm{P})/(1+\eta_{p}(\bm{P}))$ in Theorem 3.3.
One may also consider applying typical concentration inequalities for Markov sums [14, 15, 16, 20, 13, 8] to establish tail bounds similar to Theorem 2.1(2). This will require analyzing other spectral gaps of $\bm{P}_{2}$ defined in Subsection 2.3, which wil also be presented in Theorem 3.3 as a result of independent interest. Specifically, we show that $\eta_{a}(\bm{P}_{2})\equiv 0$ for any transition matrix $\bm{P}$. Hence, most typical concentration results are inapplicable to to $\bm{P}_{2}$, except for [20, Theorem 3.4]. On the other hand, we prove that $\eta_{ps}(\bm{P}_{2})\geq\eta_{ps}(\bm{P})/2$, which allows us to apply [20, Theorem 3.4] to deduce that
$\mathbb{P}\left\{{|r_{n}(u,v)-\mu(u)p(u,v)|>t}\right\}\lesssim\exp\left(-cn\eta_{ps}(\bm{P})t^{2}\right)$ (2.17)
if $n\gtrsim 1/(\eta_{ps}(\bm{P}))$. However, as shown in Subsection 2.3, if $\bm{P}$ is not geometrically ergodic, then it is possible that $\eta_{ps}(\bm{P})=0$. Moreover, since $\eta_{p}(\bm{P})\geq\eta_{ps}(\bm{P})/2$, the bound in (2.17) does not provide any improvement over the spectral gap term of Theorem 2.1(2). In fact, since all other spectral gaps introduced in Subsection 2.3 can be bounded in terms of the IP gap, the only potential benefit of employing other concentration results instead of [10] lies in achieving a possibly better constant $c$ in the resulting tail bound.
Since Theorem 2.1(2) is an entry-wise result, it is natural to ask if we can establish a probabilistic error bound for MLE at the matrix level that depends on the IP gap of $\bm{P}$. The main challenge in obtaining a result analogous to [23, Theorem 3.3] lies in studying the concentration of the error norm (2.16), since Paulin’s result [20, Corollary 2.10] requires a finite mixing time, which is not guaranteed for all irreducible transition matrices. It would be interesting to develop a Markov concentration inequality for general functions of $n$ variables instead of sums based on the IP gap. If such a result is established, one could then derive a matrix-level bound analogous to [23, Theorem 3.3]. We will leave it as a future work.
On the other hand, when $\bm{P}$ is reversible, the transition matrix $\tilde{\bm{P}}_{2}$ defined on the symmetric path space $\tilde{\Omega}^{2}$ always admits a finite mixing time, which can be shown to be bounded by the mixing time of $(\bm{P}+\bm{I})/2$ plus 1 similar to [23, Lemma 6.1]. Therefore, following the proof strategy similar to [23, Theorem 3.3], one can derive a matrix-level bound for SCE in terms of the mixing time of $(\bm{P}+\bm{I})/2$, which can further be bounded by $c\log(\mu_{*})/\eta(\bm{P})$. Here $\mu_{*}:=\min_{u\in\Omega}\mu(u)>0$ by the Perron–Frobenius theorem. We do not pursue this approach here, as the proof is essentially the same. Instead, we establish Theorem 2.2(2) by applying a recently established matrix concentration inequality [19] to $\tilde{\bm{P}}_{2}$, providing a matrix-level error bound for SCE via a more direct method. Note, however, that the result in [19] requires a nonzero absolute spectral gap of the transition matrix, which does not hold for $\bm{P_{2}}$. Therefore, extending this approach to MLE would require a matrix concentration result based on more robust spectral gaps (e.g., the IP gap), which is another interesting direction for future research.